HUBBLE SPACE TELESCOPE DAILY REPORT #1701 PERIOD COVERED: 0000Z (UTC) 08/20/96 - 0000Z (UTC) 08/21/96 Daily Status Report as of 234/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of FOS/RD 6914 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration - Red Side) The Faint Object Spectrograph (Red Side) was used to make both a = TALED and a wavelength calibration image. All observations were made = using internal light sources. This proposal will be executed a total of = 12 times during cycle 6 (i.e., once per month). The observations were = executed as scheduled, and no anomalies were noted. 1.2 Completed GHRS/1 6457 (The Abundance of Interstellar Nitrogen) The High Resolution Spectrograph (Side One) was used to make high = signal-to-noise ratio spectrographic observations of the background, OB = class star HD143275 in order to obtain absorption spectra from the = foreground interstellar nitrogen. The nitrogen abundance is important as = previous determinations of carbon and oxygen abundances show that C and O = abundances are sufficiently less than the solar values. The observations = were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 6099 (Black Holes and Cores of Galaxies) The WF/PC-2 was used to search for possible black holes in the = cores of nearby early-type galaxies, by obtaining high resolution images = of galaxy NGC-1023 to identify its bright core, and to provide = sufficiently accurate coordinates for the FOS, which will be used to = measure the velocity dispersion around this galaxy. The observations = were executed as scheduled, with no anomalies. 1.4 Completed FGS 6472 (Dynamical Masses for the Stars in the = Pre-Main-Sequence Classification) The FGS #3 was used in both the TRANS and POS modes to make = astrometric observations of the pre-main-sequence, binary system, star = STAR-04560 (aka, 045251+3016). These data will be used to measure the = dynamical mass ratio of the pair, and the inclination of their orbits. = The observations were executed as scheduled, and no anomalies were noted. = 1.5 Completed FOS/RD 6577 (An Expanded Survey for Study of the = Lyman-Alpha Line and QSO Mg II Absorption) The Faint Object Spectrograph (Red Side) was used to make = observations of the QSO PKS0421+01 in order to make a survey of the = Lyman-Alpha line and the 68 Mg II absorption line system. The = observations were completed as scheduled, with no problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: There were no reported operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: DF-224 Flight Software Version 9.7 is due to be installed = tomorrow. /DMH